OpenStudy (anonymous):

Solve. 16 + (–25) = (1 point) A. –39 B. 11 C. –9 D. 9

OpenStudy (anonymous):

-9 that is C option

OpenStudy (anonymous):

-9 C.

OpenStudy (anonymous):

+ and - becomes - hence 16+(-25) = 16-25 = -9

OpenStudy (anonymous):

C. –9